A car of mass 2500 kg is driving up a hill inclined at an angle of 10° to the horizontal. If the

resistance to motion is 3.5 kN, and the speed is a constant 20 m/s, find the power exerted.


1
Expert's answer
2021-01-04T14:39:19-0500
"v=const"


"a=0"


"F-F_r-mgsin10^0=0"


"F=F_r+mg sin10^0"


"F=3,500+2,500\times9.8\times 0.17=7,752 N"


"P=Fv"


"P=7,752\times20=155,040 W=155 kW"